OpenStudy (anonymous):

help on algebra 2 A please?

OpenStudy (destinymasha):

whats the question?

OpenStudy (anonymous):

15-12n=n+2

OpenStudy (anonymous):

13=13n n=1

OpenStudy (destinymasha):

1

OpenStudy (anonymous):

Solution for 15-12n=n+2 equation: Simplifying 15 + -12n = n + 2 Reorder the terms: 15 + -12n = 2 + n Solving 15 + -12n = 2 + n Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '-1n' to each side of the equation. 15 + -12n + -1n = 2 + n + -1n Combine like terms: -12n + -1n = -13n 15 + -13n = 2 + n + -1n Combine like terms: n + -1n = 0 15 + -13n = 2 + 0 15 + -13n = 2 Add '-15' to each side of the equation. 15 + -15 + -13n = 2 + -15 Combine like terms: 15 + -15 = 0 0 + -13n = 2 + -15 -13n = 2 + -15 Combine like terms: 2 + -15 = -13 -13n = -13 Divide each side by '-13'. n = 1 Simplifying n = 1

OpenStudy (destinymasha):

n = 1

OpenStudy (destinymasha):

you did not just write that in 10 seconds @trumpetmaster7777

OpenStudy (anonymous):

word, then copy and paste lol

OpenStudy (destinymasha):

lol, cheap way to get medals };

OpenStudy (anonymous):

thanks u so much everyone(:

OpenStudy (destinymasha):

ur welcome

OpenStudy (anonymous):

\[-14-5yge-64\]

OpenStudy (anonymous):

no, i write it on word, then copy it and paste it onto this website @Destinymasha

OpenStudy (destinymasha):

sure , look im not trying to bust you or anything, just messin around

OpenStudy (destinymasha):

lol

OpenStudy (anonymous):



OpenStudy (destinymasha):

OHHH, thats a biggie , not sure about that one

OpenStudy (anonymous):

7n-4(5-2n)=10